# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ 1.3. Main ideas
Upon Fourier transformation, the linearized operator of the first equation in (1.29) is expressed as $i\xi y-\partial_{y}^{2}$, and it shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$. Then it is straight to observe that there exists a solution $\hat{f}_{e}$ to the following linearized problem
$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big{|}_{y=0}=0,\quad\xi\in{\mathbb{R}},\end{aligned}\right$.
for any given function $h\in L^{2}(\Omega_{M})$, and $\hat{f}$ is the Fourier transformation of $f$ in the variable $x$. Moreover, similar to the proof in [9], the solution satisfies
(1.30) ${\big{\|}f_{e}\big{\|}}_{Y_{0}}\leq C{\big{\|}h\big{\|}}_{L^{2}}.$
Here $C>0$ is independent of $M$. Thus, in order to solve (1.26), we need to obtain the $L^{2}$-estimate of the nonlinear terms $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$. By calculations, we obtain
(1.31) ${\big{\|}u\partial_{x}\omega+v\partial_{y}\omega\big{\|}}_{L^{2}}\lesssim({\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}+{\big{\|}y^{\frac{1}{2}}\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}+{\big{\|}y\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}^{2}$
$\qquad\qquad\qquad\qquad\quad+{\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data}.$
Clearly, the norms ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}$ and ${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$ can not be controlled by ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}$. Even if we add low-frequency information, the subspace $L^{2}(\Omega_{M})\cap L_{x}^{2}\dot{H}_{y}^{2}\cap\dot{H}_{x}^{\frac{2}{3}}L_{y}^{2}$ of $Y_{0}$ is unable to be embedded into $L^{\infty}$. Thus, we need not only a low-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}(\Omega)}$, but also a higher-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}$ for $\alpha>0$ such that
(1.32) ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{\alpha}}^{1-\theta_{1}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{\alpha}}^{1-\theta_{1}},\ \text{for some }\theta_{1}\in(0,1),$
${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}^{\theta_{2}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}^{1-\theta_{2}},\ \text{for some }\theta_{2}\in(0,1).$
The main difficulty of this paper is to obtain the low-frequency estimate of $\bar{\omega}_{e}$. We observe that the homogeneous solution of the linearized operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$. For any nonhomogeneous term $h$, the general solution includes a term of the form $(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz$, where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function. Note that the nonlinear term $u\partial_{x}\omega-v\partial_{y}\omega$ cannot be entirely recast solely as $x$-directional derivatives (i.e., $\partial_{x}(\text{Terms})$). This structural constraint hinders low-frequency estimates derived from the nonlinear terms. To address the singularity near $\xi=0$, a key observation is that the function
@@
has no zeros in the sector $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})\}$, where $Ai(z),Bi(z)$ denote two independent Airy functions. This property is non-trivial (see Remark 3.1), which involves the distribution of zeros of the Bessel function. Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form:
(1.33) $G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
@@
It is straight to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\}$. With these observations, we derive that
(1.34) ${\big{\|}f_{e}\big{\|}}_{L^{2}}\lesssim{\big{\|}h\big{\|}}_{L^{2}}.$
To control the higher-frequency of $f_{e}$, we introduce the space $Y_{\alpha}$ for $\alpha>0$. It requires higher regularity of $h$ since
(1.35) ${\big{\|}f_{e}\big{\|}}_{Y_{\alpha}}\lesssim{\big{\|}h\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}},\quad\forall\alpha>0.$
By product estimates, we finally control the $H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$-norm of nonlinear terms in $\bar{h}$} for $\alpha>\frac{2}{3}$, that is,
${\big{\|}{\bar{h}}\big{\|}}_{H_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},\quad\forall\alpha>\frac{2}{3}.$
It follows that
(1.36) ${\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data},\quad\forall\alpha>\frac{2}{3}.$
Although the above process is sufficient to obtain a solution $(\omega,A)$ in $L^{2}\cap Y_{0}\cap Y_{\alpha}$ for $\alpha>\frac{2}{3}$, it is natural to require that the displacement $A$ is bounded for all $M\geq M_{0}$. We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then it satisfies
@@
Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ estimate of $\omega$ via the norm ${\big{\|}\omega\big{\|}}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$. To this end, we utilize the interpolation inequality. Let $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$ be an interpolation exponent, then we have
@@
It implies that the bound of $A$ can not be controlled uniformly in $M$ by the norm ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}$, which contradicts the physical observations that the linear part $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ determines the behavior of $(\omega,A)$. To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big{\|}}_{L^{2}}$, so as to derive a refined estimate. Specifically, we obtain
@@
Here $C>0$ is independent of $M$. And the estimate of the new norm is derived by using more information of Airy functions (see Proposition 2.3).
1.4. Organization of the paper
This paper is organized as follows. First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key a priori estimates, including weighted Sobolev norms for vorticity and displacement. Subsequently, Section 3 focuses on the linearized problem of the triple-deck system. By decomposing the linearized equations into well-posed subproblems (e.g., boundary-value problems for the Airy equation), we construct solutions to the linearized system and establish quantitative bounds on their regularity. Thereafter, in Section 4, we apply these linear estimates to prove the main result (Theorem 1.2) via a fixed-point iteration scheme, establishing the existence and uniqueness of solutions to the full nonlinear triple-deck equations. Finally, we present an appendix, which provides a rigorous formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer. We also present explicit solutions for the upper and main layers, demonstrating their dependence on the viscous sublayer solution. This appendix reinforces the self-contained nature of the mathematical framework.
2. Preliminaries
To study the quantity $\omega$, we first consider the space $X_{0}\cap X_{\alpha}$. Note that the space $X_{\bar{\alpha}}$ consists of a low-frequency part and a high-frequency part. The low-frequency part is characterized by the norms ${\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}$ and ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}$, while the high-frequency part is characterized by ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}$. We can explore various properties between these two parts. First, we present some of them for $f\in X_{0}$.
Proposition 2.1.
Assume that $f\in X_{0}$, for any $0\leq\varepsilon\leq\frac{1}{6}$, we have
(2.1) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}+{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}+{\big{\|}f\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big{\|}f\big{\|}}_{X_{0}}.$
Here $C$ is independent of $\varepsilon$ and $f$.
@@ 2. Preliminaries
By Sobelev embedding theorem and the interpolation theorem, it is straight to obtain
(2.2) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}H_{y}^{2}}\leq C{\big{\|}f\big{\|}}_{X_{0}},$
@@
Using the inequality ${\big{\|}|\partial_{x}|^{-\frac{1}{3}}f\big{\|}}_{L_{x}^{6}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}}$, we derive that
(2.3) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}.$
Having established the above estimates, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through some calculations, we obtain
(2.4) ${\big{\|}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg{\{}{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big{\|}y\partial_{x}f\big{\|}}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
@@
Observe that as the index $\alpha$ increases, higher frequencies in the variable $x$ is accounted for in $X_{\alpha}$, since ${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}f\big{\|}}_{Y_{0}}={\big{\|}f\big{\|}}_{Y_{\alpha}}$. These higher frequencies lead to other useful properties.
Proposition 2.2.
Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$, then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we obtain
(2.5) ${\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}+{\big{\|}f\big{\|}}_{L^{\infty}(\Omega_{M})}\leq C\bigg{(}\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\bigg{)}({\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}).$
Here $C>0$ is independent of $\varepsilon,\alpha$, and $f$.
@@
By interpolation and Proposition 2.1, we have
(2.6) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big{\|}|\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
@@
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\bigg{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\bigg{)}{\big{\|}f\big{\|}}_{Y_{0}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}({\big{\|}f\big{\|}}_{Y_{0}}+{\big{\|}f\big{\|}}_{Y_{\alpha}}).$
By classical extension, we denote $\tilde{f}$ as the extension of $f$ to ${\mathbb{R}}^{2}$. It is evident to verify that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
(2.7) ${\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
@@
By calculations, we have
(2.8) ${\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\hat{\tilde{f}}\big{\|}}_{L_{\xi,\eta}^{1}}$
@@
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}.$
Here $C(\alpha,\theta)={\big{\|}(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big{\|}}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$ in the last inequality. It implies that
(2.9) ${\big{\|}f\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big{\|}f\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big{(}{\mathbb{R}};L_{y}^{2}((0,M))\big{)}}\bigg{)}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}\bigg{)}.$
Next, we present some a priori estimates for the nonlinear terms. Suppose $(u,v)$ satisfies
$\partial_{x}u+\partial_{y}v=0,$
with the boundary data
@@
Set $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$ for $\omega=\partial_{y}u$, then we have
(2.10) $h=-u\partial_{x}\omega-\bigg{(}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz-y\partial_{x}A_{M}\bigg{)}\partial_{y}\omega.$
Define

### v2

@@ 1.3. Main difficulties and Key ideas
Taking the Fourier transformation in $x$, the linear operator of the first equation in (1.17) is expressed as $i\xi y-\partial_{y}^{2}$. It shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$. Indeed, there exists a solution $\hat{f}_{e}$ to the following linear problem
$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big|_{y=0}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
Similar to the proof in [9], the solution satisfies
${\big\|f_{e}\big\|}_{Y_{0}}\leq C{\big\|h\big\|}_{L^{2}}.$
However, for the existence of the solution, we need to establish the following $L^{2}$ boundedness
${\big\|f_{e}\big\|}_{L^{2}}\lesssim{\big\|h\big\|}_{L^{2}},$
which is the key difference and difficulty compared with the work [9]. We observe that the homogeneous solution to the linear operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$. For any inhomogeneous term $h$, the general solution includes a term of the form
$(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz,$
where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function. Note that the nonlinear term $u\partial_{x}\omega-v\partial_{y}\omega$ cannot be entirely recast solely as $x$-directional derivatives (i.e., $\partial_{x}(\text{Terms})$). This structural constraint hinders low-order estimates derived from the nonlinear terms. To address the singularity near $\xi=0$, a key observation is that the function
@@
has no zeros in the sector $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})\}$, where $Ai(z),Bi(z)$ denote two independent Airy functions. This property is non-trivial (see Remark 3.1), which involves the distribution of zeros of the Bessel function. Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form
$G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
@@
It is easy to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\big\}$.
The last key point is to show that the displacement $A$ is bounded for all $M\geq M_{0}$. We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then
@@
Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ norm of $\omega$ by means of the norm ${\big\|\omega\big\|}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$ via interpolation: there exists some $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$,
@@
This means that the bound of $A$ cannot be controlled uniformly in $M$ by the norm ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}$—a contradiction to the physical intuition that that the linear component $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ dictates the behavior of $(\omega,A)$. To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$, which allows us to derive a refined estimate. Specifically, we have
@@
Here $C>0$ is independent of $M$. The estimate for this new norm is established by leveraging additional properties of Airy functions (see Proposition 2.1).
1.4. Notations and Organization of the paper
We adopt standard mixed-norms on $\Omega_{M}$ as follows
$\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big\|f\big\|}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$ such that $B_{0}\leq CB_{1}$.
This paper is organized as follows. First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key nonlinear estimates, including weighted Sobolev norms for vorticity and displacement. Subsequently, Section 3 focuses on the solvability of the linear system of the triple-deck system. Thereafter, in Section 4, we prove the main result (Theorem 1.2) via a Picard iteration scheme. Finally, we present an appendix, which provides a formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer.
2. Functional spaces and nonlinear estimates
To study the quantity $\omega$, we first consider the space $X_{0}\cap X_{\alpha}$. Note that the space $X_{\bar{\alpha}}$ consists of a low-order part and a high-order part. The low-order part is characterized by the norms ${\big\|f\big\|}_{L^{2}(\Omega_{M})}$ and ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}$, while the high-order part is characterized by ${\big\|f\big\|}_{Y_{\bar{\alpha}}}$. We now present some properties between these two parts for $f\in X_{0}$.
Lemma 2.1.
It holds that for any $0\leq\varepsilon\leq\frac{1}{6}$,
${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}+{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}+{\big\|f\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big\|f\big\|}_{X_{0}}.$
Here $C$ is independent of $\varepsilon$.
@@ 2. Functional spaces and nonlinear estimates
By Sobelev embedding theorem and the interpolation, we obtain
${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big\|f\big\|}_{L_{x}^{2}H_{y}^{2}}\leq C{\big\|f\big\|}_{X_{0}},$
@@
Using the inequality ${\big\||\partial_{x}|^{-\frac{1}{3}}f\big\|}_{L_{x}^{6}}\leq C{\big\|f\big\|}_{L_{x}^{2}}$, we derive
${\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}.$
Having established the estimates above, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through direct calculations, we obtain
${\big\|f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg\{{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big\|y\partial_{x}f\big\|}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
@@
Lemma 2.2.
Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$. Then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we have
${\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{\infty}}+{\big\|f\big\|}_{L^{\infty}(\Omega_{M})}\leq C\Big(\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\Big)\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
Here $C>0$ is independent of $\varepsilon,\alpha$.
@@
By the interpolation and Lemma 2.1, we have
${\big\|\partial_{y}f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big\||\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
@@
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\Big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big\||\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\Big){\big\|f\big\|}_{Y_{0}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\big({\big\|f\big\|}_{Y_{0}}+{\big\|f\big\|}_{Y_{\alpha}}\big).$
We denote by $\tilde{f}$ the extension of $f$ to ${\mathbb{R}}^{2}$ such that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
${\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
@@
Direct calculations show that
${\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\hat{\tilde{f}}\big\|}_{L_{\xi,\eta}^{1}}$
@@
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big).$
Here $C(\alpha,\theta)={\big\|(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big\|}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$. This shows that
${\big\|f\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big)$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big\|f\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big({\mathbb{R}};L_{y}^{2}((0,M))\big)}\big)$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
Next we are devoted to the estimates of nonlinear terms. Let the velocity $(u,v)$ satisfy the divergence-free condition
$\partial_{x}u+\partial_{y}v=0$
subject to the boundary conditions
@@
where $A_{M}$ denotes a fixed velocity at the upper boundary $y=M$. Define the function
$h=-(u\partial_{x}\omega+v\partial_{y}\omega),$
for $\omega=\partial_{y}u$. Using the divergence-free condition and boundary conditions, we can rewrite $h$ as
$h=-u\partial_{x}\omega-(V^{0}-y\partial_{x}A_{M})\partial_{y}\omega.$
Here